40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1418     Entry time: Tue May 17 21:53:36 2011
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: new excess noise issue resolved 

[Alastair, Zach]

We have been trying to figure out why we see (more) excess low frequency noise in the current configuration. The ~100x optical gain imbalance could have been an unrelated issue, but it seemed like a reasonable place to start today.

We began by bypassing the PDs altogether and just injecting part of the LO signal directly into the RF input of the mixers (where the RFPD signal comes in). This should create a DC signal at the low-passed IF output, which can be maximized by tuning the LO phases with the dials on the boxes. Using equal input RF levels, the maximized outputs of the respective mixers were different by a factor of ~100 (220 mV to 2 mV). We double-checked that the LO levels were equal, and since they were we deduced that something was wrong with the primary box's demod electronics. Checking each of the parts individually, we found that the ZX05 mixer was busted. We replaced it and the optical gain is now 100x higher. This explains the discrepancy we have been noticing from the theoretically achievable level for months.

I reduced the gain of the servo to accommodate the optical gain, but we still needed >10dB of attenuation between the PD and the mixer to keep the loop from oscillating. This needs investigation.

When we plugged everything back in and took a spectrum, we found that the low-frequency noise was back to the level (a bit lower, in fact) where it was before we reduced the optical power from 30 mW to 3 mW. It is a bit worse at higher frequencies, but this is the best we have ever done in the current configuration. Plot:

back_to_normal.png

It is a bit disheartening that there isn't a real improvement, but we must consider this:

  • We have increased the primary loop optical gain by a factor of 100. This cannot ever be a bad thing.
  • We need to redo the measurements from the last post to verify that electronics noise isn't the problem.
  • If electronics noise is an issue, we need to redistribute gain to take advantage of the higher optical gain.
  • It's not obvious that the old and new spectra have the same shape in the requirement band. This means that it's quite possible we have removed the old limiting noise source by reducing clipping on the FIs, and that some other non-ideality is limiting us right now. 
ELOG V3.1.3-